# Generative Adversarial Networks for High-Dimensional Item Factor Analysis: A Deep Adversarial Learning Algorithm

### v1

@@ 3.2 Adaptive Contrast Technique
$\max_{\bm{\theta}}\max_{\bm{\phi}}\;\mathbb{E}_{p_{\bm{D}}(\bm{x})}\left[-\text{KL}\left[q_{\bm{\phi}}(\bm{z}\mid\bm{x}),r_{\alpha}(\bm{z}\mid\bm{x})\right]+\mathbb{E}_{q_{\bm{\phi}}(\bm{z}\mid\bm{x})}\left[-\log r_{\alpha}(\bm{z}\mid\bm{x})+\log p_{\bm{\theta}}(\bm{x},\bm{z})\right]\right].\$ (23)
@@ 3.3 Importance-weighted Technique
$\bm{z}_{1:R}$ $\sim\prod\limits_{r=1}^{R}q_{\bm{\phi}}(\bm{z}_{r}\mid\bm{x})$
$w_{r}$ $=p_{\bm{\theta}}(\bm{z}_{r},\bm{x})/q_{\bm{\phi}}(\bm{z}_{r}\mid\bm{x})$
@@ 3.4 Implementation Summary
Given : Responses $\bm{x}_{1},\dots,\bm{x}_{N}$; dimension of latent variable $P$; mini-batch size $B$; IW samples $R$; MC samples $S$; hyperparameters for AdamW $\beta_{1}$ and $\beta_{2}$; learning rate for the encoder and decoder $\eta_{1}$; learning rate for the discriminator $\eta_{2}$; initialization parameters $\bm{\theta}_{0}$, $\bm{\phi}_{0}$ and $\bm{\psi}_{0}$; prior distribution $p(\bm{z})$ (e.g., multivariate standard Normal distribution)
for iteration $t=1$ to $T$ do
@@
       $\bm{m}_{t}=\beta_{1}\bm{m}_{t-1}+(1-\beta_{1})\bm{g}_{t}$;
       $\bm{v}_{t}=\beta_{2}\bm{v}_{t-1}+(1-\beta_{2})\bm{g}_{t}^{2}$;
       $\bm{g}_{t+1}=\bm{g}_{t}-\eta\frac{\bm{m}_{t}}{\sqrt{\bm{v}_{t}}+\epsilon}-\eta\lambda\bm{g}_{t}$;
@@
The learning rate in the AdamW gradient update significantly impacts the model’s efficiency and effectiveness, as this hyperparameter influences the speed at which the model converges to an optimal solution. A learning rate that is too small may result in exceedingly slow learning or convergence to an undesired local minimum, while a learning rate that is too large can cause oscillations between different local minima. To address these challenges, cyclical learning rate (CLR) schedules (Smith,, 2017) have been developed to dynamically adjust the learning rate throughout the training process. Unlike traditional fixed or monotonically decreasing learning rates, CLR oscillates the learning rate within predefined bounds, enabling the model to escape local minima and explore the optima more effectively. To ensure that the discriminator can learn more rapidly and provide more informative gradients to the generator, the discriminator can be assigned a higher learning rate compared to the generator (Heusel et al.,, 2017). The base values for the CLR were selected from the range $\{1,2,3,4,5,6,7,8,9,10\}\times 10^{-3}$ for the discriminator and $\{1,2,3,4,5,6,7,8,9,10\}\times 10^{-4}$ for the encoder. The upper bound for the CLR was set to five times the base value. Additionally, experiments were conducted to identify the optimal base values for different tasks.
To ensure the convergence of the algorithm, the IW-ELBO value for each mini-batch is recorded. Every 100 fitting iterations, the average IW-ELBO across these 100 mini-batches is calculated. The fitting process is terminated, and convergence is considered achieved if this average does not improve after 500 such calculations.
In exploratory studies, although different optimization runs may produce varied estimates, some results can be shown to be equivalent after suitable transformation. The equivalence is checked by comparing the loadings matrices from each run, following the process suggested by Urban and Bauer, (2021). First, the factors are rotated using the Geomin oblique rotation method (Yates,, 1987). Next, any factor whose loadings sum to a negative value is reversed (Asparouhov and Muthén,, 2009) so that all factors point in the same direction. Then, one matrix is chosen as a reference while the columns in the other matrices are rearranged to reduce the mean-squared error between matching entries. After aligning the matrices, Tucker’s congruence coefficient is calculated (Lorenzo-Seva and Ten Berge,, 2006); if this value is above 0.98, the solutions are considered equivalent (MacCallum et al.,, 1999). The same process of inversion and rearrangement is applied to factor correlation matrices, where both rows and columns are reordered for a direct comparison.
@@ 4.2 Simulation Study: Latent Variables Following Normal Distribution
The true latent variables were generated as 500 samples from a multivariate normal distribution with a zero mean and the true correlation matrix. For the item parameters, each item’s true loading value was sampled from a log-normal distribution, obtained by exponentiating samples from $\mathcal{N}(0,0.5)$. The true intercept parameters were generated from another multivariate normal distribution with a zero mean and a covariance matrix drawn from the LKJ distribution, then sorted into strictly increasing order. Using the conditional probabilities defined by Equation (2), a categorical distribution was employed to sample 100 different response matrices. Both the IWAVB and IWAE methods were fitted to these simulated response datasets and compared in terms of estimation accuracy, computation time, and approximate log-likelihood. The estimation accuracy of the true latent variables or parameters, $\xi$, was assessed using the mean-square error (MSE):
@@
Figure 8 shows box plots of the MSEs for item parameters (loadings and intercepts), latent variable values, and factor correlations for the IWAVB and IWAE methods with 15, 25, or 35 importance-weighted (IW) samples. Both methods achieved low MSEs for estimating item loadings and factor correlations. While the MSE for item intercepts was relatively higher, it still averaged below 0.03, which is satisfactory. The MSE for latent variable estimates was the highest due to the large number of respondents but did not exhibit extreme values. Although the IWAE method slightly outperformed IWAVB in estimating item intercepts, latent variables, and factor correlations, the differences were minimal. Notably, IWAVB achieved a smaller MSE for item loadings. Both methods showed marginal improvements in estimation accuracy as the number of IW samples increased.
@@ 4.3 Simulation Study: Latent Variables Following Multimodal Distribution
This simulation study aimed to evaluate a more general case where the latent variables follow a multimodal distribution. A graded response model was configured with $P=1$ latent factor, $N=500$ respondents, $M=20$ items, and $C_{j}=3$ response categories for $j=1,\dots,M=20$. The true latent variables were sampled from a mixture of three Gaussian distributions: $\mathcal{N}(-1.5,0.5)$, $\mathcal{N}(0,0.5)$, and $\mathcal{N}(1.5,0.5)$, with respective weights of $0.4$, $0.2$, and $0.4$, as shown in Figure 10. This resulted in a distribution with multiple modes, in contrast to the unimodal structure of a normal distribution. For the item parameters, 20 true loading values were generated from a log-normal distribution with a mean of zero and variance of $0.5$. True intercepts were sampled using a similar procedure as described in the previous simulation study.
@@
We use the following base learning rates: $\eta_{1}=0.003$ for the discriminator and $\eta_{2}=0.0002$ for the encoder in the IWAVB method, and $\eta=0.0005$ for the inference model (one hidden layer with 100 neurons) in the IWAE method. The configuration of neural networks was same as the previous study. The number of IW samples was set to $R=25$ for both methods. Using these configurations, Figure 11 presents box plots comparing the MSEs for item loadings, item intercepts, and latent variable values estimated by the IWAVB and IWAE methods. The IWAVB method consistently achieved lower MSEs across all parameter estimates compared to IWAE.
@@
Figure 12: Computation time and approximate log-likelihood comparison for IWAVB and IWAE methods across 100 replications of simulation with latent variables following multimodal distribution.

### v2

@@ 3.2 Adaptive Contrast Technique
$\max_{\bm{\theta}}\max_{\bm{\phi}}\;\mathbb{E}_{p_{\bm{D}}(\bm{x})}\left[-\text{KL}\left[q_{\bm{\phi}}(\bm{z}\mid\bm{x}),r_{\alpha}(\bm{z}\mid\bm{x})\right]+\mathbb{E}_{q_{\bm{\phi}}(\bm{z}\mid\bm{x})}\left[-\log r_{\alpha}(\bm{z}\mid\bm{x})+\log p_{\bm{\theta}}(\bm{x},\bm{z})\right]\right].$ (22)
@@ 3.3 Importance-weighted Technique
$\bm{z}_{1:R}$ $\sim\prod\limits_{r=1}^{R}q_{\bm{\phi}}(\bm{z}_{r}\mid\bm{x}),w_{r}=p_{\bm{\theta}}(\bm{z}_{r},\bm{x})/q_{\bm{\phi}}(\bm{z}_{r}\mid\bm{x})$
@@ 3.4 Implementation Summary
Given : Responses $\bm{x}_{1},\dots,\bm{x}_{N}$; dimension of latent variable $P$; mini-batch size $B$; IW samples $R$; MC samples $S$; hyperparameters for AdamW $\beta_{1}$ and $\beta_{2}$; learning rate for the encoder and decoder $\eta_{1}$; learning rate for the discriminator $\eta_{2}$; initialization parameters $(\bm{\theta}_{0},\bm{\phi}_{0},\bm{\psi}_{0})$ with Kaiming and Xavier Initialization (preventing issues like vanishing or exploding gradients); prior distribution $p(\bm{z})$ (e.g., multivariate standard Normal distribution)
for iteration $t=0$ to $T$ do
@@
       $\tilde{\bm{g}}_{t}=\bm{g}_{t}+\lambda(\bm{\theta}_{t}^{T},\bm{\phi}_{t}^{T},\bm{\psi}_{t}^{T})$, $\bm{m}_{t+1}=\beta_{1}\bm{m}_{t}+(1-\beta_{1})\tilde{\bm{g}}_{t}$, $\bm{v}_{t+1}=\beta_{2}\bm{v}_{t}+(1-\beta_{2})\tilde{\bm{g}}_{t}^{2}$;
       $\hat{\bm{m}}_{t+1}=\bm{m}_{t+1}/(1-\beta_{1}^{t+1})$, $\hat{\bm{v}}_{t+1}=\bm{v}_{t+1}/(1-\beta_{2}^{t+1})$;
       $(\bm{\theta}_{t+1}^{T},\bm{\phi}_{t+1}^{T},\bm{\psi}_{t+1}^{T})=(\bm{\theta}_{t}^{T},\bm{\phi}_{t}^{T},\bm{\psi}_{t}^{T})-\eta\frac{\hat{\bm{m}}_{t+1}}{\sqrt{\hat{\bm{v}}_{t+1}}+\epsilon}-\eta\lambda(\bm{\theta}_{t}^{T},\bm{\phi}_{t}^{T},\bm{\psi}_{t}^{T})$;
@@
The learning rate in the AdamW gradient update significantly impacts the model’s efficiency and effectiveness, as this hyperparameter affects the speed at which the model converges to an optimal solution. A learning rate that is too small may result in exceedingly slow learning or convergence to an undesired local minimum, while a learning rate that is too large can cause oscillations between different local minima. To address these challenges, cyclical learning rate (CLR) schedules (Smith,, 2017) have been developed to dynamically adjust the learning rate throughout the training process. Unlike traditional fixed or monotonically decreasing learning rates, CLR oscillates the learning rate within predefined bounds, enabling the model to escape local minima and explore the optima more effectively. To ensure that the discriminator can learn more rapidly and provide more informative gradients to the generator, the discriminator can be assigned a higher learning rate compared to the generator (Heusel et al.,, 2017). The base values for the CLR were selected from the range $\{1,2,3,4,5,6,7,8,9,10\}\times 10^{-3}$ for the discriminator and $\{1,2,3,4,5,6,7,8,9,10\}\times 10^{-4}$ for the encoder, and the upper bound was set to five times the base value. Experiments were conducted to identify the optimal base values for different tasks.
To ensure the convergence of the algorithm, the IW-ELBO value for each mini-batch is recorded. Every 100 fitting iterations, the average IW-ELBO across these 100 mini-batches is calculated. If this average does not improve after 500 such calculations, the fitting process is terminated, and convergence is considered achieved.
In exploratory studies, although different optimization runs may produce varied estimates, some results can be shown to be equivalent after suitable transformation. The equivalence is checked by comparing the loading matrices from each run, following the process suggested by Urban and Bauer, (2021). First, the factors are rotated using the Geomin oblique rotation method (Yates,, 1987). Next, any factor whose loadings sum to a negative value is reversed (Asparouhov and Muthén,, 2009) so that all factors point in the same direction. Then, one matrix is chosen as a reference while the columns in the other matrices are rearranged to reduce the mean-squared error between matching entries. After aligning the matrices, Tucker’s congruence coefficient is calculated (Lorenzo-Seva and Ten Berge,, 2006); if this value is above 0.98, the solutions are considered equivalent (MacCallum et al.,, 1999). The same process of inversion and rearrangement is applied to factor correlation matrices, where both rows and columns are reordered for a direct comparison.
@@ 4.2 Simulation Study: Latent Variables Following Normal Distribution
The true latent variables were generated as 500 samples from a multivariate normal distribution with a zero mean and the true correlation matrix. For the item parameters, each item’s true loading value was sampled from a log-normal distribution, obtained by exponentiating samples from $\mathcal{N}(0,0.5)$. The true intercept parameters were generated from another multivariate normal distribution with a zero mean and a covariance matrix drawn from the LKJ distribution, then sorted into strictly increasing order. Using the conditional probabilities defined by Equation (2), a categorical distribution was employed to sample 100 different response matrices. Both the IWAVB and IWAE methods were fitted to these simulated response datasets and compared in terms of estimation accuracy, computation time, and approximate log-likelihood. The estimation accuracy of the true parameters, $\xi$, was assessed using the mean-square error (MSE):
@@
Figure 8 shows box plots of the MSEs for item parameters (loadings and intercepts), and factor correlations for the IWAVB and IWAE methods with 15, 25, or 35 importance-weighted (IW) samples. Both methods achieved low MSEs for estimating item loadings and factor correlations. While the MSE for item intercepts was relatively higher, it still averaged below 0.03, which is satisfactory. Although the IWAE method slightly outperformed IWAVB in estimating item intercepts and factor correlations, the differences were minimal. Notably, IWAVB achieved a smaller MSE for item loadings. Both methods showed marginal improvements in estimation accuracy as the number of IW samples increased.
@@ 4.3 Simulation Study: Latent Variables Following Multimodal Distribution
This simulation study aimed to evaluate a more general case where the latent variables follow a multimodal distribution. A graded response model was configured with $P=1$ latent factor, $N=500$ respondents, $M=20$ items, and $C_{j}=3$ response categories for $j=1,\dots,M=20$. The true latent variables were sampled from a mixture of three Gaussian distributions: $\mathcal{N}(-1.5,0.5)$, $\mathcal{N}(0,0.5)$, and $\mathcal{N}(1.5,0.5)$, with respective weights of $0.4$, $0.2$, and $0.4$, as shown in Figure 10. This resulted in a distribution with multiple modes, in contrast to the unimodal structure of a normal distribution. For the item parameters, 20 true loading values were generated from a log-normal distribution with a mean of zero and variance of $0.5$. True intercepts were sampled using a similar procedure as described in the previous simulation study. Additionally, a three-dimensional case was examined by configuring a GRM with $P=3$ latent factor, $N=5000$ respondents, $M=45$ items, and $C_{j}=3$ per item. Each item also corresponded to only one factor. In this instance, the true latent variables were generated from a mixture of three trivariate Gaussian distributions with mean vectors of the form $(\mu,\mu,\mu)$, where $\mu$ takes on the values $-1.5$, 00, and $1.5$, and with a fixed correlation matrix sampled from the LKJ distribution. The true loading and intercept values were generated using analogous procedures as described previously.
@@
We used the following base learning rates for the discriminator and encoder in the IWAVB method: $(\eta_{1}=0.003,\eta_{2}=0.0002)$ for $P=1$ and $(\eta_{1}=0.002,\eta_{2}=0.0005)$ for $P=3$, while for the inference model (one hidden layer with 100 neurons) in the IWAE method, both cases utilized $\eta=0.0005$. The configuration of neural networks was same as the previous study. The number of IW samples was set to $R=30$ for both methods. Using these configurations, Figure 11 presents box plots comparing the MSEs for item loadings and intercepts estimated by the IWAVB and IWAE methods. The IWAVB method consistently achieved lower MSEs across both parameter estimates compared to IWAE.
@@
Figure 12: Approximate log-likelihood and computation time comparison for IWAVB and IWAE methods across 100 replications of simulation with latent variables following multimodal distribution.